A farmhouse shelters 11 animals. Some are goats and some are ducks. Altogether there are 34 legs. How many of each animal are there?
Answer by ikleyn(52879)

.
A farmhouse shelters 11 animals. Some are goats and some are ducks.
Altogether there are 34 legs. How many of each animal are there?
~~~~~~~~~~~~~


Let x be the number of goats.

Then the number of ducks is  11-x.


There is no need to remind that a goat has 4 legs, while a duck has only 2.


Write this simple equation counting the legs


    4x + 2*(11-x) = 34.


Simplify and solve


    4x + 22 - 2x = 34

    4x - 2x = 34 - 22

      2x    =    12

       x    = 12/2 = 6.


ANSWER.  6 goats and  11-6 = 5 ducks.

Solved.

-----------------

To see many other similar solved problems, look into the lessons
    - Problem on two-wheel and three-wheel bicycles
    - Problem on animals at a farm
    - Problem on pills in containers
    - What type of problems are these?
in this site.